I am in a College Algebra class online and could use a little help with a word problem regarding banking.
The question is:
Wendy, a loan officer at a bank, has 1,000,000 to lend and is required to obtain an average return of 18% per year. If she can lend at the rate of 19% or at the rate of 16%, how much can she lend at the rate of 16% and still meet her requirement?
I am thinking that you would use the formula I=PRT but I am unsure of how to set up this problem.

Answer by Alan3354(69443)

Wendy, a loan officer at a bank, has 1,000,000 to lend and is required to obtain an average return of 18% per year. If she can lend at the rate of 19% or at the rate of 16%, how much can she lend at the rate of 16% and still meet her requirement?
--------------
18% of 1000000 = 180000 she has to make on interest.
---
s = amount @ 16%
--
0.16s + 0.19(1000000 - s) = 180000
16s + 19000000 - 19s = 18000000
-3s = -10000000
s = $333,333.33